answersLogoWhite

0

🎒

Statistics

Statistics deals with collecting, organizing, and interpreting numerical data. An important aspect of statistics is the analysis of population characteristics inferred from sampling.

36,755 Questions

How does PCIe Version 2 compare to Version 1?

PCIe version 2 doubled the fequency of the PCIe bus, theoretically doubling the throughput, It also allows for up to 32 lanes on one slot

If using numbers 1 thru 49 how many 6 digit combinations can be made?

Using the formula n!/r!(n-r)! where n is the number of possible numbers and r is the number of numbers chosen, there are 13983816 combinations of six numbers between 1 and 49 inclusive.

How to calculate the null and alternative hypothesis and test 5 level of significance?

The null and alternative hypotheses are not calculated. They should be determined before any data analyses are carried out.

What are the benefits of cumulative frequency?

It shows what proportion of the total population are less than (or equal to) each value.

What is the outlier in this sequence 51014151720354850?

One number, such as 51014151720354850, does not form a sequence.

What is the 75 rounded nearest 10 is?

The answer can be 70 or 80 since 75 is equally far away from both. Generally you pick an answer appropriate to the context, for example; if you need to sit 75 people for a meal and the tables seat 10 each, then you will want to round up to 80.

The rules of rounding require 5 to round up. The answer is 80.

If I have all combinations for my state cash 5 lottery game is there a strategy to figure out which numbers will be drawn next?

No. If there were any way to figure out which numbers will be drawn next,

it would be a fixed, rigged, dishonest game.

When two dice are rolled find the probability of not rolling a sum of 5?

When rolling a pair of dice there are 6∙6 = 36 possible outcomes.

The outcomes that give a sum of 5 are 4, [(1,4), (2,3), (3,2), (4,1)].

So the probability of not rolling a sum of 5 is:

P(NOT 5) = 1 - P(5) = 1 - 4/36 = 32/36 = 8/9 = 0.8888... ≈ 88.9%

What z value corresponds to a confidnece level of 86?

The value for a one-sided confidence interval of 86% is 1.08

What is the voluntary code of practice?

Voluntary codes are codes of practice that can also be called codes of conduct or voluntary initiatives. The non-legislatively required commitments are designed to shape, influence, benchmark, or control behavior. One or more organization or person agrees to them and they are to be applied in consistent manners to attain consistent outcomes.

What does this 0x00e5415b referenced memory 0x00000000 error mean?

If you got this while your screen turned blue and the computer restarted, then it could mean a problem with your computer hardware. I would start by checking your ram with memtest. Also if you are running an overclock, remove it and run at stock settings.

There are 3 step to repair memory 0x00000000 error

If you got memory 0x00000000 error then there is a 94% chance that your computer has registry problems. To repair memory 0x00000000 error you need to follow the steps below:

* Step 1 - Download a Perfect Optimizer,install this error repair tool.

* Step 2 - Click the Repair All Button.It will scan you PC for Free.

* Step 3 - Then click the Repair All Button again and your done! It is very easy to repair memory 0x00000000 error.

What does hither mean?

As an adverb "hither" means "to or toward this place". As an adjective it means "located on the near side."